OpenStudy (anonymous):

I really would appreciate some help: If f(x)=[(x^2)sin(1/x), x not = 0, c, x=0] c is a constant, Fill in the blanks with polynomials: _____≤(x^2)sin(1/x)≤_____ I have no idea how to even begin, please help

OpenStudy (jamesj):

\[ | \sin x | \leq 1 \] hence \[ | \sin(1/x) | \leq 1 \]

OpenStudy (anonymous):

would you mind explaining that a little more, sorry about the trouble

OpenStudy (jamesj):

sin(x) oscillates between -1 and 1. So we can bound it below by -1 and above by +1 In fact, for any function f, sin(f(x)) still only varies between -1 and 1. Given that, you should be able to see now how to bound your expression above and below

OpenStudy (jamesj):

To be explicit, for all \( x \neq 0 \), \[ -1 \leq \sin(1/x) \leq 1 \]

OpenStudy (jamesj):

Here's the plot

OpenStudy (anonymous):

I see but the problem asks for polynomials I think that is what is confusing me, how would you find them?

OpenStudy (jamesj):

\( x^2 \) is a polynomial

OpenStudy (anonymous):

then the blanks would be -x^2 and x^2?

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

I think i have a better idea now, thank you for your help and patience!

OpenStudy (jamesj):

here's another example, \[ -x^2 \leq \sin(e^{-x^2} + x^{100}) x^2 \leq x^2 \] This is true because sin can't be any more than +1 or any less than -1, despite whatever the form of the argument of sin(.) is.

OpenStudy (anonymous):

I understand, its because of the sin then, once again thank you!